a circuit consisting of 9 volt battery and two 4ohm resistors connected in series. How many Coulomb flows through each resistor in a second?


1
Expert's answer
2021-06-21T16:27:27-0400

Gives

Battery (V)=9V

"R_1=R_2=4\Omega"

Both are series combination

"R_{eq}=R_1+R_2=4+4=8\Omega"

"V=IR"

"I=\frac{V}{R}=\frac{9}{8}=1.125A"

Each resistance flow charg(coulamb )

"Q=It=1\times1.125=1.125C"